# Optimization of Excitation Waveforms for Maximum Instantaneous Field Intensity

### v2

@@ V Conclusion
The capability of the tool is to find the optimal excitation waveform for an arbitrarily chosen target metric. Moreover, the number of sources is not restricted. In the case of an antenna, matching can be omitted. This leads to fundamental bounds based on current density as known in the frequency domain, but now the current, apart from space, can be controlled in time allowing us to evaluate fundamental bounds in the time domain for a given design region based on the current, although the values would in many cases be overly optimistic.
An important strength of the proposed method lies in its practical utility for evaluating and optimizing radiating systems. Given a specific system configuration, the method determines whether the desired pulsed performance can be achieved. If the shape and constraints of the system are sufficient, the method identifies the optimal pulse shape to maximize performance. Conversely, if the system cannot meet the imposed requirements, the approach highlights these limitations, offering valuable insights into whether a redesign of the system geometry is necessary.
@@
By decoupling waveform optimization from structural design, the proposed method ensures a globally optimal solution that can be computed efficiently in polynomial time, unlike the iterative and often local solutions associated with topology optimization. This characteristic guarantees the general applicability of the method to existing antenna systems without necessitating their redesign, while still ensuring the best achievable performance for a given scenario and set of constraints.
@@
While this paper focuses on waveform optimization for a fixed transmitting antenna, the methodology has broader potential. It can be extended to co-design scenarios in which both antenna and waveform design are considered simultaneously, making it applicable to a wider range of advanced engineering problems. In such case, the topology optimization would change the design according to the results computed by the proposed tool, though such a procedure would be computationally intensive. However, this would open the way to the automated design of pulsed radiators, antiferromagnetic memories, or other devices that rely on pulses in the time domain.
@@ Appendix A Matrix Operators from EFIE and MoM
A representative example of an EM solver capable of providing required antenna characteristics is AToM based on EFIE and MoM. The matrix representation of multiple quantities is presented in [39, Appendix A]. The extraction of essential metrics to the manuscript is presented in this appendix. Multiple ports are considered.
@@
The essential quantities are far field as a function of the incident waves from the feeding transmission line and their total available energy in (4). The electric far field is computed from far-field matrix $\mathbf{F}$ and current vector $\mathbf{I}$, see (13). The current is excited by incident waves $\mathbf{a}$ as
@@
Appendix B Band Limited Set of Basis Functions
@@
Appendix C Projection via Galerkin Method
Galerkin method [38, 13] converts continuous operators to matrices using the same basis and testing functions. The method is applied to all operators using functions introduced in Appendix B. The approach is explicitly presented on the far-field operator and the construction of the far-field power matrix.
The far-filed operator is projected in the following way for optimized time $t_{\mathrm{opt}}$
$\hat{\mathbf{F}}_{pn}(\boldsymbol{r},\omega)=\\ \int\limits_{-\infty}^{\infty}\mathbf{F}\left(\boldsymbol{r},\omega\right)\mathbf{Z}^{-1}\left(\omega\right)\mathbf{P}\mathbf{k}_{\mathrm{i}}^{-1}\left(\omega\right)\xi_{n}\left(\omega\right)\mathrm{e}^{\mathrm{j}\omega t_{\mathrm{opt}}}\,\mathrm{d}\omega,$ (17)
where $\mathbf{a}_{p}$ excites only the chosen feeding transmission line $p$ by unity amplitude. Here, it is presented for operators from MoM, see Appendix A, but the same methodology can be applied to antenna characteristics from an arbitrary EM solver. The far-field operator projected has two rows corresponding to polarizations and $pn$ columns with a linear index that combines ports and basis functions. The far-filed operator allows us to define the far-field power matrix
$\mathbf{S}(\boldsymbol{r},\omega)=\hat{\mathbf{F}}^{\mathrm{H}}(\boldsymbol{r},\omega)\hat{\mathbf{F}}(\boldsymbol{r},\omega)$ (18)
in (8).

### v3

@@ VIII Conclusion
A methodology to determine optimal waveforms is presented in this paper with optimized parameters being defined as linear or quadratic terms. The evaluation requires a post-processing step based on convex optimization and the input data can be calculated in any EM simulator. The technique can serve and provide quantitative results in many scenarios, including pulsed radiation or antiferromagnetic memory switching with pulse lasers, as discussed in this paper, as well as the creation of hotspots using near-field energy focusing in lossy materials using antenna arrays for applications such as hyperthermia therapy in medicine. The capability of the tool is to find the optimal excitation waveform for an arbitrarily chosen target metric. Moreover, the number of sources is not restricted. In the case of an antenna, matching can be omitted. This leads to fundamental bounds based on current density as known in the frequency domain, but now the current, apart from space, can be controlled in time allowing us to evaluate fundamental bounds in the time domain for a given design region based on the current, although the values would in many cases be overly optimistic.
@@
An important strength of the proposed method lies in its practical utility for evaluating and optimizing radiating systems. Given a specific system configuration, the method determines whether the desired pulsed performance can be achieved. If the shape and constraints of the system are sufficient, the method identifies the optimal pulse shape to maximize performance. Conversely, if the system cannot meet the imposed requirements, the approach highlights these limitations, offering valuable insights into whether a redesign of the system geometry is necessary. By decoupling waveform optimization from structural design, the proposed method ensures a globally optimal solution that can be computed efficiently in polynomial time, unlike the iterative and often local solutions associated with topology optimization. This characteristic guarantees the general applicability of the method to existing systems without necessitating their redesign, while still ensuring the best achievable performance for a given scenario and set of constraints.
@@
The methodology has broader potential. It is extended to co-design scenarios in which both antenna and waveform design are considered simultaneously, making it applicable to a wider range of advanced engineering problems. In such case, the topology optimization changes the design according to the results computed by the proposed tool. The combination of the pulse and topology design is present on an end-fire two-dipole array, resulting in a significant improvement of the performance.
@@ Appendix A Matrix Operators from EFIE and MoM
A representative example of an EM solver capable of providing required antenna characteristics is AToM based on EFIE and MoM. The matrix representation of multiple quantities is presented in [30, Appendix A]. The extraction of essential metrics to the manuscript is presented in this appendix. Multiple ports are considered.
@@
The essential quantities are far field as a function of the incident waves from the feeding transmission line and their total available energy. The electric far field is computed from far-field matrix $\mathbf{F}$ and current vector $\mathbf{I}$, see (12). The current is excited by incident waves $\mathbf{a}$ as
@@
Appendix B Energy and Field Operators
Fourier transform is used to transform target quantities from the time domain to the frequency domain. The mapping applies to energy and field expressions.
The field operator for a given location or direction, and time reads
$\boldsymbol{E}(\boldsymbol{r},t)=\\ \int\limits_{-\infty}^{\infty}\mathbf{H}\left(\boldsymbol{r},\omega\right)\mathbf{a}\left(\omega\right)\mathrm{e}^{\mathrm{j}\omega t}\,\mathrm{d}\omega.$ (15)
One of the common targets is intensity $\left|\boldsymbol{E}\left(\boldsymbol{r},t\right)\right|^{2}$ at optimized time $t_{\mathrm{opt}}$. Choosing an arbitrary point in the space $\boldsymbol{r}$ and time window of length $2T$, another important measure is energy delivered to that point in that time window which reads
$W(\boldsymbol{r},T)=C\int\limits_{t_{\mathrm{d}}-T}^{t_{\mathrm{d}}+T}\left|\boldsymbol{E}\left(\boldsymbol{r},t\right)\right|^{2}\mathrm{d}t=\\ C\int\limits_{\Omega}\int\limits_{\Omega}\dfrac{2\sin\left[\left(\omega_{1}-\omega_{2}\right)T\right]}{\omega_{1}-\omega_{2}}\mathrm{e}^{-\mathrm{j}\left(\omega_{1}-\omega_{2}\right)t_{\mathrm{d}}(\boldsymbol{r})}\\ \mathbf{a}^{\mathrm{H}}\left(\omega_{2}\right)\mathbf{H}^{\mathrm{H}}(\boldsymbol{r},\omega_{2})\mathbf{H}(\boldsymbol{r},\omega_{1})\mathbf{a}\left(\omega_{1}\right)\mathrm{d}\omega_{2}\mathrm{d}\omega_{1},$ (16)
where $\Omega:\omega\in[-\omega_{\mathrm{max}},-\omega_{\mathrm{min}}]\cup[\omega_{\mathrm{min}},\omega_{\mathrm{max}}]$ is the intgration domain, $C$ is a constant to fit the units and $t_{\mathrm{d}}$ is the time delay from the excitation source.
For incident energy, the transfer function $\mathbf{H}$ is omitted and the delay time is zero
$W_{\mathrm{in}}(\boldsymbol{r},T)=C\int\limits_{t_{\mathrm{d}}-T}^{t_{\mathrm{d}}+T}\left|\mathbf{a}\left(t\right)\right|^{2}\mathrm{d}t=\\ C\int\limits_{\Omega}\int\limits_{\Omega}\dfrac{2\sin\left[\left(\omega_{1}-\omega_{2}\right)T\right]}{\omega_{1}-\omega_{2}}\mathbf{a}^{\mathrm{H}}\left(\omega_{2}\right)\mathbf{a}\left(\omega_{1}\right)\mathrm{d}\omega_{2}\mathrm{d}\omega_{1}.$ (17)
The expression is simplified for the total energies, which read,
$W(\boldsymbol{r},T)=C\int\limits_{-\infty}^{\infty}\left|\boldsymbol{E}\left(\boldsymbol{r},t\right)\right|^{2}\mathrm{d}t=\\ C\int\limits_{\Omega}\mathrm{e}^{-\mathrm{j}\omega t_{\mathrm{d}}(\boldsymbol{r})}\mathbf{a}^{\mathrm{H}}\left(\omega\right)\mathbf{H}^{\mathrm{H}}(\boldsymbol{r},\omega)\mathbf{H}(\boldsymbol{r},\omega)\mathbf{a}\left(\omega\right)\mathrm{d}\omega,$ (18)
or
$W_{\mathrm{in}}(\boldsymbol{r},T)=C\int\limits_{-\infty}^{\infty}\left|\mathbf{a}\left(t\right)\right|^{2}\mathrm{d}t=C\int\limits_{\Omega}\mathbf{a}^{\mathrm{H}}\left(\omega\right)\mathbf{a}\left(\omega\right)\mathrm{d}\omega,$ (19)
Galerkin method [27, 28] converts continuous operators shown above to matrices using the same basis and testing functions. The method is applied to all operators using functions introduced in Appendix C.
Appendix C Band Limited Set of Basis Functions